# Combining GCN Structural Learning with LLM Chemical Knowledge for or Enhanced Virtual Screening

### v1

@@
Combining GCN Structural Learning with LLM Chemical Knowledge for or Enhanced Virtual Screening
@@ 1 Abstract
Virtual screening plays a critical role in modern drug discovery by enabling the identification of promising candidate molecules for experimental validation. Traditional machine learning methods such as support vector machines (SVM) and XGBoost rely on predefined molecular representations, often leading to information loss and potential bias. In contrast, deep learning approaches—particularly Graph Convolutional Networks (GCNs)—offer a more expressive and unbiased alternative by operating directly on molecular graphs. Meanwhile, Large Language Models (LLMs) have recently demonstrated state-of-the-art performance in drug design, thanks to their capacity to capture complex chemical patterns from large-scale data via attention mechanisms.
@@ 3.1 Descriptor-based methods
These approaches rely on engineered features derived from molecular fingerprints, physicochemical properties, and atom-level descriptors, which are calculated using cheminformatics toolkits. Open-source tools such as RDKit are widely used for generating these descriptors due to their accessibility and flexibility, making them a popular choice in academic and industrial settings. However, for more advanced and specialized descriptors, commercial software such as Dragon and MOE (Molecular Operating Environment) is often required, which can be expensive and introduce additional complexity to the workflow. Despite the effectiveness of descriptor-based models such as Support Vector Machines (SVMs) [34, 22] and XGBoost [3], their reliance on these features makes the process time-consuming, particularly for large datasets, and heavily dependent on the quality of the calculated descriptors. These limitations highlight the need for more scalable and automated approaches to molecular representation in drug design. However, these methods are often limited by the quality and completeness of the handcrafted features, making them less effective for capturing the complex structural relationships inherent in molecules. Handcrafted descriptors often fail to account for non-linear interactions and higher-order dependencies between molecular substructures, which are critical for modeling intricate molecular activities. For example, descriptors such as molecular fingerprints may encode local structural features but overlook long-range interactions or stereochemical properties, which are essential for accurate property prediction and activity classification [27, 28]. Moreover, the manual selection and design of features can introduce human bias, reducing the generalizability of these methods to novel datasets or diverse chemical spaces [31].
@@ 3.2 Graph representation approaches
To overcome the limitations of descriptor-based methods, which rely on predefined features and may fail to capture the full complexity of molecular structures, graph representation methods particularly Graph Convolutional Networks (GCNs) [18] have emerged as a powerful alternative to encode the cervical cell [29] . These approaches model molecules as graphs, where nodes represent atoms and edges represent bonds, allowing for a more natural and flexible representation of molecular structures. Unlike handcrafted descriptors, graph-based methods can dynamically learn molecular features directly from the graph topology, capturing both local and global structural relationships. This capability addresses key shortcomings of descriptor-based approaches, such as their inability to model non-linear interactions and long-range dependencies between molecular substructures. GCNs have been successfully applied to molecular property prediction tasks such as predicting the inhibitors of the 5 CYP isoforms [24]. Despite their ability to learn structural features, traditional GCNs excel at capturing the relational and topological information inherent in molecular graphs. By representing molecules as graphs, where nodes correspond to atoms and edges represent bonds, GCNs can model both local substructures and their connectivity within the molecular framework[30, 1]. This approach enables GCNs to extract detailed structural information, such as bond order, aromaticity, and atom hybridization, making them particularly effective for tasks that rely heavily on understanding molecular topology, such as property prediction and activity classification. Furthermore, graph representations inherently encode spatial relationships between molecular components, which are challenging to capture using descriptor-based or text-based methods alone.
@@ 3.3 SMILES approaches
Recently, SMILES (Simplified Molecular Input Line Entry System), a widely used textual representation of molecular structures [33], has gained prominence due to its compatibility with natural language processing techniques. The ability to leverage sequence-based models has facilitated the encoding of molecular structures into meaningful numerical representations.
@@ 4 Methodology
The proposed approach introduces a novel integration of Graph Convolutional Networks (GCNs) with SMILES encodings derived from a pretrained large language model (LLM), bridging the gap between structural and semantic molecular representations. In this architecture, the molecular graph provides a detailed account of the relational and topological features, such as element symbol, formal charge and neighbors, while the SMILES encodings capture rich chemical semantics, including functional group interactions and stereochemistry. By embedding the SMILES information directly into each layer of GCN layers, the model dynamically incorporates global chemical knowledge learned from extensive molecular corpora during the LLM’s pretraining phase. To facilitate this integration, a projection layer was employed to transform the SMILES embedding into a 10-dimensional space before fusion with the learned graph representations. This transformation ensures that molecular descriptors are properly aligned with the GCN’s feature space, allowing the model to better capture interactions between structural and semantic information. This dual-stream approach enables the model to remember the global features calculated by the LLM trained on large number of smiles and overcome the limitations of traditional GCNs, which rely solely on graph-based features, and descriptor-based methods, which depend on handcrafted features that are often incomplete or computationally expensive to generate.
This synergistic design allows the model to learn a comprehensive and context-aware representation of molecular properties, enhancing its predictive power across diverse datasets. The integration of LLM-derived SMILES embeddings in each layers of GCN not only enriches the feature space but also provides the model with a broader understanding of chemical space, enabling it to generalize more effectively to unseen molecules. This capability positions the proposed GCN-LLM hybrid as a powerful tool for advancing molecular property prediction and drug discovery tasks, addressing key challenges in cheminformatics and demonstrating significant potential for real-world applications.
@@ 4.4.1 Evaluation Metrics
For the sake of comparison between SVM, XGBoost, GCN, and our hybrid model (GCN-LLM), we carefully preprocessed all datasets. This preprocessing involved a stratified sampling approach to ensure a balanced distribution of classes within the training and test sets. The performance of the models was assessed using a comprehensive set of metrics, including accuracy, F1-score, precision, recall, and the area under the ROC curve (AUC-ROC). These metrics were chosen to provide a nuanced understanding of the models’ capabilities, capturing both overall classification accuracy and their ability to distinguish between positive and negative instances. This multifaceted evaluation approach enables a thorough analysis of each model’s strengths and weaknesses, offering valuable insights into their performance across different aspects of the molecular classification task. Therefore, we incorporated the following metrics:
@@ 4.4.4 Area Under the ROC Curve (AUC-ROC)
By using this multifaceted evaluation framework, we validate not only the overall predictive power of the models but also their ability to manage class imbalances and correctly identify molecular properties critical to downstream tasks. This thorough analysis ensures a nuanced understanding of the strengths and limitations of each approach, providing valuable insights into their suitability for diverse molecular classification tasks.
@@ 5 Results and Discussion
As demonstrated in Table 1 and figures, the GCN-LLM model consistently outperformed all baseline methods across multiple datasets, achieving the highest accuracy (91.2%) and F1-score (90.8%) on Dataset erbB1. This remarkable performance highlights the model’s ability to seamlessly integrate both structural and semantic features of molecular data, a critical factor in improving classification accuracy for complex molecular systems. The integration of SMILES embeddings generated by a pretrained large language model (LLM) adds a new dimension to molecular representation, capturing intricate chemical semantics that are often overlooked by traditional descriptor-based methods. Unlike SVM and XGBoost, which rely on handcrafted molecular descriptors, the GCN-LLM dynamically incorporates semantic-rich SMILES features into each graph convolutional layer. This enables the model to refine node-level embeddings iteratively, leveraging both the structural properties of molecular graphs and the chemical context provided by SMILES strings.

### v2

# Combining GCN Structural Learning with LLM Chemical Knowledge for Enhanced Virtual Screening

@@
Combining GCN Structural Learning with LLM Chemical Knowledge for Enhanced Virtual Screening
@@ 1 Abstract
Virtual screening plays a critical role in modern drug discovery by enabling the identification of promising candidate molecules for experimental validation. Traditional machine learning methods such, as Support Vector Machines (SVM) and XGBoost, rely on predefined molecular representations, often leading to information loss and potential bias. In contrast, deep learning approaches—particularly Graph Convolutional Networks (GCNs)—offer a more expressive and unbiased alternative by operating directly on molecular graphs. Meanwhile, Large Language Models (LLMs) have recently demonstrated state-of-the-art performance in drug design, thanks to their capacity to capture complex chemical patterns from large-scale data via attention mechanisms.
@@ 3.1 Descriptor-based methods
These approaches rely on engineered features derived from molecular fingerprints, physicochemical properties, and atom-level descriptors, which are calculated using cheminformatics toolkits. Open-source tools such as RDKit are widely used for generating these descriptors due to their accessibility and flexibility, making them a popular choice in academic and industrial settings. However, for more advanced and specialized descriptors, commercial software such as Dragon and MOE (Molecular Operating Environment) is often required, which can be expensive and introduce additional complexity to the workflow. Despite the effectiveness of descriptor-based models such as Support Vector Machines (SVMs) [35, 23] and XGBoost [3], their reliance on these features makes the process time-consuming, particularly for large datasets, and heavily dependent on the quality of the calculated descriptors. These limitations highlight the need for more scalable and automated molecular representation methods in drug design, such as graph-based and SMILES-based approaches. However, these methods are often limited by the quality and completeness of the manual feature design, making them less effective for capturing the complex structural relationships inherent in molecules. Handcrafted descriptors often fail to account for non-linear interactions and higher-order dependencies between molecular substructures, which are critical for modeling intricate molecular activities. For example, descriptors such as molecular fingerprints may encode local structural features but overlook long-range interactions or stereochemical properties, which are essential for accurate property prediction and activity classification [28, 29]. Moreover, the manual selection and design of features can introduce human bias, reducing the generalizability of these methods to novel datasets or diverse chemical spaces [32].
@@ 3.2 Graph representation approaches
To overcome the limitations of descriptor-based methods, which rely on predefined features and may fail to capture the full complexity of molecular structures, graph representation methods particularly Graph Convolutional Networks (GCNs) [18] have emerged as a powerful alternative for encoding molecular structures [30]. These approaches model molecules as graphs, where nodes represent atoms and edges represent bonds, allowing for a more natural and flexible representation of molecular structures. Unlike handcrafted descriptors, graph-based methods can dynamically learn molecular features directly from the graph topology, capturing both local and global structural relationships. This capability addresses key shortcomings of descriptor-based approaches, such as their inability to model non-linear interactions and long-range dependencies between molecular substructures. GCNs have been successfully applied to molecular property prediction tasks such as predicting the inhibitors of the 5 CYP isoforms [25]. Despite their ability to learn structural features, traditional GCNs excel at capturing the relational and topological information inherent in molecular graphs. By representing molecules as graphs, where nodes correspond to atoms and edges represent bonds, GCNs can model both local substructures and their connectivity within the molecular framework[31, 1]. This approach enables GCNs to extract detailed structural information, such as bond order, aromaticity, and atom hybridization, making them particularly effective for tasks that rely heavily on understanding molecular topology, such as property prediction and activity classification. Furthermore, graph representations inherently encode spatial relationships between molecular components, which are challenging to capture using descriptor-based or text-based methods alone.
@@ 3.3 SMILES approaches
Recently, SMILES (Simplified Molecular Input Line Entry System), a widely used textual representation of molecular structures [34], has gained prominence due to its compatibility with natural language processing techniques. The ability to leverage sequence-based models have facilitated the encoding of molecular structures into meaningful numerical representations.
@@ 4 Methodology
The proposed approach introduces a novel integration of Graph Convolutional Networks (GCNs) with SMILES encodings derived from a pretrained large language model (LLM), bridging the gap between structural and semantic molecular representations. In this architecture, the molecular graph provides a detailed account of the relational and topological features, such as element symbol, formal charge and neighbors, while the SMILES encodings capture rich chemical semantics, including functional group interactions and stereochemistry. By embedding the SMILES information directly into each layer of GCN layers, the model dynamically incorporates global chemical knowledge learned from extensive molecular corpora during the LLM’s pretraining phase. To facilitate this integration, a projection layer was employed to transform the SMILES embedding into a 10-dimensional space before fusion with the learned graph representations. This transformation ensures that the SMILES embeddings are properly aligned with the GCN’s feature space, facilitating the model’s ability to capture interactions between structural and semantic information. This dual-stream approach enables the model to remember the global features calculated by the LLM trained on large number of SMILES and overcome the limitations of traditional GCNs, which rely solely on graph-based features, and descriptor-based methods, which depend on handcrafted features that are often incomplete or computationally expensive to generate.
This synergistic design allows the model to learn a comprehensive and context-aware representation of molecular properties, enhancing its predictive power across diverse datasets. The integration of LLM-derived SMILES embeddings in each layers of GCN not only enriches the feature space but also provides the model with a broader understanding of chemical space, enabling it to generalize more effectively to unseen molecules. This makes the proposed GCN-LLM hybrid a powerful tool for advancing molecular property prediction and drug discovery tasks, addressing key challenges in cheminformatics and demonstrating significant potential for real-world applications.
@@ 4.4.1 Evaluation Metrics
For the sake of comparison between SVM, XGBoost, GCN, and our hybrid model (GCN-LLM), we carefully preprocessed all datasets. This preprocessing involved a stratified sampling approach to ensure a balanced distribution of classes within the training and test sets. The performance of the models was assessed using a comprehensive set of metrics, including accuracy, F1-score, precision, recall, and the area under the ROC curve (AUC-ROC). These metrics were chosen to provide a detailed understanding of the models’ capabilities, capturing both overall classification accuracy and their ability to distinguish between positive and negative instances. This multifaceted evaluation approach enables a thorough analysis of each model’s strengths and weaknesses, offering valuable insights into their performance across different aspects of the molecular classification task. Therefore, we incorporated the following metrics:
@@ 4.4.4 Area Under the ROC Curve (AUC-ROC)
By using this multifaceted evaluation framework, we validate not only the overall predictive power of the models but also their ability to manage class imbalances and correctly identify molecular properties critical to downstream tasks. This thorough analysis ensures a comprehensive understanding of the strengths and limitations of each approach, providing valuable insights into their suitability for diverse molecular classification tasks.
@@ 5 Results and Discussion
As demonstrated in Table 4 and Figures 2 and 3, the GCN-LLM model consistently outperformed all baseline methods across multiple datasets, achieving the highest accuracy (91.2%) and F1-score (90.8%) on the erbB1 dataset. This remarkable performance highlights the model’s ability to seamlessly integrate both structural and semantic features of molecular data, a critical factor in improving classification accuracy for complex molecular systems. The integration of SMILES embeddings generated by a pretrained large language model (LLM) adds a new dimension to molecular representation, capturing intricate chemical semantics that are often overlooked by traditional descriptor-based methods. Unlike SVM and XGBoost, which rely on handcrafted molecular descriptors, the GCN-LLM dynamically incorporates semantic-rich SMILES features into each graph convolutional layer. This enables the model to refine node-level embeddings iteratively, leveraging both the structural properties of molecular graphs and the chemical context provided by SMILES strings.